OpenStudy (zyberg):

Solve system of equations for rational roots: x^2 + y^4 = 20 x^4 + y^2 = 20. I have tried quite a lot of things, but I can't seem to grasp the right path.

OpenStudy (caozeyuan):

you can redefine your variables

OpenStudy (caozeyuan):

define x^2=m, y^2=n

OpenStudy (caozeyuan):

m+n^2=20, m^2+n=20

OpenStudy (caozeyuan):

now you subtract, so (m-n)+(n^2-m^2)=0

OpenStudy (caozeyuan):

so (m-n)+(n+m)(n-m)=0

OpenStudy (caozeyuan):

now, if n is not the same as m, it does not have rational solutions ( plug it into a calculator and see)

OpenStudy (caozeyuan):

so m=n is the only solution

OpenStudy (caozeyuan):

so m is either -5 or 4, but we cant have imaginary number solutions so m=4 and x=y=+or - 2

OpenStudy (zyberg):

"plug it into a calculator". What if I don't have one at hand, when solving?

OpenStudy (caozeyuan):

you first need to get to the equation (m-n)+(n+m)(n-m)=0,by change of variable

OpenStudy (caozeyuan):

then you recognize that there are two cases to consider, 1) n=m and 2)n=\=m, the first one is easy, you say that m+m^2=20 and solve using calculator or quratic formula

OpenStudy (caozeyuan):

the second part is alittle bit tricky, if n is not m, then n-m is not zero, so divide by n-m on both sides of the new equation you just derived

OpenStudy (caozeyuan):

which gives you -1+n+m=1 or n+m=1

OpenStudy (caozeyuan):

you than plug it into the equation m+n^2=20, and solve for either n or m

OpenStudy (caozeyuan):

and the answer should be irrational

OpenStudy (caozeyuan):

which you discard and call it a failure, so you have only 4 answer x=+2 or -2 and y=+2 or -2, this is 4 pairs of soution